# Statistical learning theory and Occam’s razor: The argument from empirical risk minimization

### v1

@@ 3.2. VC dimension as a measure of simplicity
In contrast to definitions of the complexity of an individual hypothesis, definitions of the capacity of a hypothesis class (like VC dimension) do not depend on a specific representation or parametrization, and do therefore possess a certain objectivity or robustness. [^22] More precisely, the capacity of a hypothesis class does not depend on how the individual hypotheses are described: all that matters is their data coverage. Language-relativity only turns up when we start redescribing the instance space (cf. Steel, 2009, p. 482). But does VC dimension also give an objective or robust measure of the simplicity of a hypothesis class? One might deny this on exactly the grounds that VC dimension does not necessarily align with natural parametrizations of individual hypotheses (Domingos, 1999, p. 413). In the case of the usual parametrization of polynomials, the higher the number of free parameters, the higher the capacity of the corresponding hypothesis class; but in other cases the two can come apart. The standard example is the class of sine curves $\{h_{\alpha}\}_{\alpha\in\mathbb{R}}$ with $h_{\alpha}(x)=\sin\alpha x$ (ibid.; Harman and Kulkarni, 2007, pp. 72f). The elements in this class are given by only one parameter (and in that sense the function class is very simple), yet the class has infinite VC dimension (Vapnik, 2000, p. 78). [^23] Vapnik (1998, p. 698) himself writes that since Occam’s razor says that the explanation with “the smallest number of features (free parameters)” is best, and since this is not supported by theoretical results, “Occam’s razor principle is misleading and perhaps should be discarded in the statistical theory of inference” (ibid., p. 699). Also see Cherkassky and Mulier (2007, pp. 146ff).
@@
Of course, this objection relies on some claim that usual parameterizations do (and exclusively do) track simplicity. But even aside from the ultimate non-robustness of representational notions of complexity, there simply exist different and sometimes conflicting intuitions. From one way of looking at it, the class of sine functions is maximally complex: exactly because every possibly data configuration is covered by it (cf. Romeijn, 2017). This is the intuition of richness or complexity (or also falsifiability [^24] Vapnik (2000, pp. 42ff) links the capacity of hypothesis classes to Popper’s falsifiability of theories. Popper famously equated simplicity with falsifiability, and introduced a quantitative notion of falsifiability of theories that he claimed aligned with number of free parameters. Corfield et al. (2009) and Harman and Kulkarni (2007, p. 50ff) argue that VC dimension is a better measure of falsifiability, though these authors appear to resist then also linking VC dimension to simplicity. ) made precise in a capacity measure. VC dimension is not the measure, but it is a plausible and robust measure of the complexity of a hypothesis class. This is enough for my purpose: if a justification is to be had for preferring low or finite capacity, then I think it is reasonable to call this a justification for preferring simplicity—even if there are other reasonable conceptions of simplicity, and even if (to stress again) this notion of simplicity pertains to hypotheses classes and not to individual hypotheses.
@@ 4. The justification for simplicity
The fundamental theorem of statistical learning theory ties the simplicity—the VC dimension—of a hypothesis class to its learnability, and indeed already to its learnability by the ERM method. This result offers, first of all, a certain justification for the ERM procedure; although this is a justification with several qualifications, chief among them its model-relativity (section 4.1). Nevertheless, I will argue that the model-relative justification that learning theory can offer fits right in with a plausible broader epistemological perspective on machine learning methods (section 4.2). Finally, I will argue that this justification for the method of ERM leads to a qualified justification for a simplicity preference (section 4.3).
@@ 4.1. The justification for empirical risk minimization
The fundamental theorem shows that ${}_{\mathcal{H}}$, for VC class $\mathcal{H}$, is a good method. It is good, and good epistemically, because it satisfies a guarantee of attaining a certain epistemic goal. This guarantee therefore constitutes an epistemic justification for the method.
@@ 4.1.1. Qualifications
A first elementary point is that the fundamental theorem is a mathematical result. Any epistemic justification derived from it, in the context of a (real-world) learning problem, needs a story how it maps to this learning problem. Most obviously, for any particular real-world problem, a meaningful application of the fundamental theorem (and justificatory claims derived from it) depends on how well the problem can be modelled in the statistical learning theory framework. This includes the match of our prior assumptions with the formal assumption of i.i.d. sampling of data from some unknown distribution, but also the match of our goals with the formal choice of the 0/1 error function. The representation of a learning problem in the formal framework of statistical learning theory already forces us to commit to and encode various assumptions (von Luxburg and Schölkopf, 2011, pp. 683f), and anything that follows from the mathematics should ideally be appraised with an eye to whether these made sense for the original learning problem.
@@
One possible reservation is that these guarantees are quintessential frequentist guarantees. They say something about what we can expect with high probability before the sampling and the learning. It is in that sense that we can call methods satisfying these guarantees reliable (Harman and Kulkarni, 2007). But these guarantees (unlike Bayesian ones) do not strictly say anything about what we can infer after the learning—what we can infer about the hypothesis that has actually been selected, other than that it has been selected by a reliable method (von Luxburg and Schölkopf, 2011, pp. 699f). For instance, the “wysiwyg” guarantee of uniform convergence does not strictly say anything about what we have gotten when we see the result. It is easy to misinterpret such guarantees.
@@
A second possible reservation is that the guarantees are rather weak. In particular, since the guarantees are agnostic about the true distribution, they are worst-case bounds that for many real-world situations—where we feel we can exclude certain classes of (“pathological”) distributions—would be overly loose or pessimistic (von Luxburg and Schölkopf, 2011, p. 680, pp. 683f). [^25] That sample sizes can in practice be much smaller was already shown by early experimental results (Cohn and Tesauro, 1992). This motivates, for instance, the study of “fast rates” under further assumptions on the distributions [^26] Or more precisely, on the relation between the distribution, hypothesis class, and loss function (see van Erven et al., 2015). and of guarantees that hold for all distributions but are still distribution-dependent. [^27] This is the idea of the “theory of universal learning” of Bousquet et al. (2021). Their motivation is that the usual “distribution-independent definition of learnability is too pessimistic to explain practical machine learning,” showing the need for alternatives that “better capture the practice of machine learning, but still give rise to a canonical mathematical theory of learning rates” (ibid., p. 533). These studies yield a more complicated picture of the conditions for and the nature of the justification for ERM. Still, I think that the fundamental theorem gives at least a plausible theoretical justification for the method of ERM. This does not exclude that one might (also) wish for different kinds of justification, in any specific problem or in general. In any case, my aim here is to flag the above qualifications, to make clear that accepting the justification for ERM, and indeed the justification for simplicity to follow, presupposes accepting those qualifications. One can reject my below argument simply by rejecting the presuppositions of the statistical learning theory framework. For instance, the reasons Kelly (2008, 2011) offers for rejecting arguments for Occam’s razor from statistical learning theory are essentially a rejection of the predictive framework. [^28] Kelly (2008, p. 329) writes that Occam’s razor “should help one to select the true theory from among the alternatives,” whereas arguments based on risk minimization do not concern “theoretical truth” but “passive prediction” (ibid., 335). “But beliefs are for guiding action and actions can alter the world so that the sampling distribution we drew our conclusions from is altered as well” (ibid.); moreover, “it is clear that the over-fitting story depends, essentially, upon noise in the data […] One would prefer that the connection between simplicity and theoretical truth not depend essentially upon randomness” (ibid.). These points are all well-taken (in particular the problem of distribution-shift has recently received more attention, Wiles et al., 2022), but are all already concerns about the scope of statistical learning theory itself. But, I will argue, conditional on those qualifications—including the presuppositions of the framework—we can formulate a justificatory argument.
@@
There is, however, a final aspect to this justification for ERM that requires more discussion. Namely, this justification can be seen to cover only one side of an inevitable trade-off.
@@ 4.1.2. The bias-complexity trade-off
The opposing pull of these two error terms is also referred to as the bias-complexity trade-off. A lower complexity—lower capacity—class excludes more possibilities, and as such embodies, in machine learning terminology, a stronger inductive bias. [^29] This is a bit more general than the bias-variance trade-off (see Hastie et al., 2009, sects. 2.6, 2.9). Complexity and (inductive) bias are here in the first instance used as informal terms (even if complexity can be made precise as VC dimension, and some authors refer to the approximation error itself as the bias), while the bias and variance in the latter are well-defined statistical terms in regression with mean squared error. The fundamental theorem yields a guarantee about finding the best in a given class, but this is inevitably a class with some inductive bias. The resulting justification must therefore be relative to this class or inductive bias. [^30] Of course, the hypothesis class is not the only—or even the most important (von Luxburg and Schölkopf, 2011, p. 684)—way in which assumptions or biases enter: as discussed in sect. 4.1 above, important modeling assumptions must already be made in the formalization of the learning problem (including choice of feature space and loss function). But discussions of inductive bias (in particular around the no-free-lunch theorems introduced shortly) normally assume that these elements of the learning problem are already in place, and center on the further inductive assumptions required.
@@
In the terminology of Sterkenburg and Grünwald (2021), this is a model-relative justification. It is indeed a justification for a learning method, ERM, that is explicitly model-dependent. The method of ERM is a very general procedure—a “generic learning rule” (Shalev-Shwartz and Ben-David, 2014, p. 68)—yet one that must, by definition, on each application be supplied with further, domain-specific, assumptions. One can view ERM as instantiating a two-place function, that apart from a data sample, also takes a particular hypothesis class. On each specific application, it must be supplied with a hypothesis class or inductive model that constitutes further context-dependent assumptions, the inductive bias. Correspondingly, the learnability guarantee for ERM is model-relative, because the notion of learnability is relative to a given inductive model or VC hypothesis class.
@@
The analysis of Sterkenburg and Grünwald aims to explain how general learning-theoretic guarantees for generic algorithms are consistent with the skeptical import of the so-called no-free-lunch theorems of supervised learning (going back to Wolpert, 1992, 1996; Schaffer, 1994). Modern versions of these results (Shalev-Shwartz and Ben-David, 2014, pp. 36ff; Sterkenburg and Grünwald, 2021, pp. 9990f) say that there can exist no universal learning algorithm: every particular algorithm is inadequate—in an absolute sense—in some possible learning situations, situations where another algorithm is adequate. And since there can be no a priori justification for privileging particular learning situations, so the further interpretation goes, there can be no theoretical justification for any particular algorithm.
@@
However, Sterkenburg and Grünwald observe that such results presuppose a conception of learning algorithms as purely data-driven or data-only: as instantiating a one-place function that only takes data samples. Thus, rather than the generic yet model-dependent ERM algorithm, the no-free-lunch statement applies to any particular instantiation of ERM with an inductive model $\mathcal{H}$, any particular one-place function ${}_{\mathcal{H}}$. The no-free-lunch result of Shalev-Shwartz and Ben-David (2014, thrm. 5.1) essentially states that for any specific inductive model $\mathcal{H}$, the data-only algorithm ${}_{\mathcal{H}}$ is inadequate (i.e., with high probability suffers high absolute error) for some situations (i.e., for some true distributions; informally, those that do not match $\mathcal{H}$’s inductive bias) where another data-only algorithm, like ${}_{\mathcal{H^{\prime}}}$ for another inductive model $\mathcal{H}^{\prime}$ (that does match the situation), is adequate. [^31] Another way of casting this result is that the class of all classifiers is not learnable: since, for any possible distribution, this class has minimum approximation error, its learnability (guarantee of low estimation error) would guarantee low absolute error for any possible distribution. In fact, the proof of Shalev-Shwartz and Ben-David (2014, thrm. 5.1) already shows the failure of learnability of classes with infinite VC dimension, and the no-free-lunch theorem is in their presentation part of the proof of the fundamental theorem (ibid., pp. 45ff).
@@
In other words, while we can have a model-relative justification for model-dependent algorithms (of the rough form, “for any instantiated inductive model of the right form, works well relative to the model”), there is no absolute justification (“works well whenever”) for any data-only algorithm. The impossibility of such absolute justification is also an important part of the argument of Domingos (1998, 1999) against a possible justification of Occam’s razor.
@@ 4.1.4. The failure of an absolute justification for simplicity
Domingos takes issue with what he calls the “second razor”: the idea that “given two [hypotheses] with the same [empirical] error, the simpler one should be preferred because it is likely to have lower generalization error” (1999, p. 410). (To be dinstinguished from the “first razor,” that “the simpler [hypothesis] should be preferred because simplicity is desirable in itself,” ibid.) [^32] We are here, of course, likewise concerned with an epistemic justification for Occam’s razor, not with claims that simplicity is better because it points to hypotheses or theories that are easier to work with or more aesthetically pleasing (cf. Sober, 2015, pp. 58f). He writes that theoretical “zero-sum arguments”—the no-free-lunch theorems and variants—“imply that the second razor cannot be true” (1999, p. 413). Namely, “they imply that, for every domain where a simpler [hypothesis] is more accurate than a more complex one, there exists a domain where the reverse is true, and thus no argument which is preferable in general can be made” (ibid.).
This is true, and an expression of the impossibility of absolute justification, applied to the choice of a simple single hypothesis. [^33] This point has also been brought out experimentally (e.g., Schaffer, 1993; Webb, 1996). However, I already observed that the relevant notion of simplicity attaches to hypothesis classes, not single hypotheses. Domingos’s critique of the theoretical “PAC-learning argument” for the second razor, that is in essence also a no-free-lunch observation, is more relevant for our purposes:
@@
4.2. The epistemology of machine learning methods
4.2.1. “Forwards-looking” epistemology

### v2

@@ 3.2. VC dimension as a measure of simplicity
In contrast to definitions of the complexity of an individual hypothesis, definitions of the capacity of a hypothesis class (like VC dimension) do not depend on a specific representation or parametrization, and do therefore possess a certain objectivity or robustness. [^22] More precisely, the capacity of a hypothesis class does not depend on how the individual hypotheses are described: all that matters is their data coverage. Language-relativity only turns up when we start redescribing the instance space (cf. Steel, 2009, p. 482). But does VC dimension also give an objective or robust measure of the simplicity of a hypothesis class?
One might deny this on exactly the grounds that VC dimension does not necessarily align with natural parametrizations of individual hypotheses (Domingos, 1999, p. 413). In the case of the usual parametrization of polynomials, the higher the number of free parameters, the higher the capacity of the corresponding hypothesis class; but in other cases the two can come apart. The standard example is the class of sine curves $\{h_{\alpha}\}_{\alpha\in\mathbb{R}}$ with $h_{\alpha}(x)=\sin\alpha x$ (ibid.). The elements in this class are given by only one parameter (and in that sense the function class is very simple), yet the class has infinite VC dimension (Vapnik, 2000, p. 78). [^23] Vapnik (1998, p. 698) himself writes that since Occam’s razor says that the explanation with “the smallest number of features (free parameters)” is best, and since this is not supported by theoretical results, “Occam’s razor principle is misleading and perhaps should be discarded in the statistical theory of inference” (ibid., p. 699). Also see Cherkassky and Mulier (2007, pp. 146ff).
@@
Of course, this objection relies on some claim that usual parameterizations do (and exclusively do) track simplicity. But even aside from the ultimate non-robustness of representational notions of complexity, there just exist different and sometimes conflicting intuitions. From one way of looking at it, the class of sine functions is maximally complex: exactly because so many possible data configuration can be fit by it (cf. Romeijn, 2017). This is the intuition of richness or complexity (or also falsifiability [^24] Vapnik (2000, pp. 42ff) links the capacity of hypothesis classes to Popper’s falsifiability of theories. Popper famously equated simplicity with falsifiability, and introduced a quantitative notion of falsifiability of theories that he claimed aligned with number of free parameters. Corfield et al. (2009) and Harman and Kulkarni (2007, p. 50ff) argue that VC dimension is a better measure of falsifiability, though these authors appear to resist linking VC dimension to simplicity. ) made precise in a capacity measure. VC dimension is not the measure, but it is a plausible and robust measure of the complexity of a hypothesis class. This is enough for my purpose: if a justification is to be had for preferring low capacity, then I think it is reasonable to call this a justification for preferring simplicity—even if there are other reasonable conceptions of simplicity, and even if (to stress again) this notion of simplicity pertains to hypotheses classes and not to individual hypotheses.
@@ 4. The justification for simplicity
The fundamental theorem of statistical learning theory ties the simplicity—the VC dimension—of a hypothesis class to its learnability, and indeed already to its learnability by the ERM method. This result offers, first of all, a certain justification for the ERM procedure; although this is a justification with several qualifications, chief among them its model-relativity (section 4.1). Nevertheless, I will argue that the model-relative justification that learning theory can offer fits right in with a plausible broader epistemological perspective on machine learning methods (section 4.2). Finally, I will assemble from all of the previous elements a qualified justification for a simplicity preference (section 4.3).
@@ 4.1. The justification for empirical risk minimization
The fundamental theorem shows that ERMH, for VC class $\mathcal{H}$, is a good method. It is good, and good epistemically, because it satisfies a guarantee of attaining a certain epistemic goal. This guarantee therefore constitutes an epistemic justification for the method.
@@ 4.1.1. Qualifications
A first elementary point is that the fundamental theorem is a mathematical result. Any epistemic justification derived from it, in the context of a real-world learning problem, needs a story how it maps to this learning problem. Most obviously, for any particular real-world problem, a meaningful application of the fundamental theorem (and justificatory claims derived from it) depends on how well the problem can be modelled in the statistical learning theory framework. This includes the match of our prior assumptions with the formal assumption of i.i.d. sampling of data from some unknown distribution, but also the match of our goals with the formal choice of the 0/1 error function. The representation of a learning problem in the formal framework of statistical learning theory already forces us to commit to and codify various assumptions (von Luxburg and Schölkopf, 2011, pp. 683f), and anything that follows from the mathematics should be appraised with an eye to whether these made sense for the original learning problem.
@@
One possible reservation is that these guarantees are quintessential frequentist guarantees. They say something about what we can expect with high probability before the sampling and the learning. In that sense we can call methods satisfying these guarantees reliable (Harman and Kulkarni, 2007). But these guarantees do not strictly say anything about what we can infer after the learning—about the hypothesis that has actually been selected, other than that it has been selected by a reliable method (von Luxburg and Schölkopf, 2011, pp. 699f). For instance, the “wysiwyg” guarantee of uniform convergence does not strictly say anything about what we have gotten when we see the result. It is easy to misinterpret such guarantees.
@@
A second possible reservation is that the guarantees may be overly weak. In particular, since the guarantees are agnostic about the true distribution, they are worst-case bounds that for many real-world situations—where we feel we can exclude certain classes of (“pathological”) distributions—would be overly loose or pessimistic (von Luxburg and Schölkopf, 2011, p. 680, pp. 683f). [^25] That sample sizes can in practice be much smaller was already shown by early experimental results (Cohn and Tesauro, 1992). This motivates, for instance, the study of “fast rates” under further assumptions on the distributions [^26] Or more precisely, on the relation between the distribution, hypothesis class, and loss function (see van Erven et al., 2015). and of guarantees that hold for all distributions but are still distribution-dependent. [^27] This is the idea of the “theory of universal learning” of Bousquet et al. (2021). Their motivation is that the usual “distribution-independent definition of learnability is too pessimistic to explain practical machine learning,” showing the need for alternatives that “better capture the practice of machine learning, but still give rise to a canonical mathematical theory of learning rates” (ibid., p. 533). These studies yield a more complicated picture of the justification for ERM.
Still, the fundamental theorem gives at least a plausible theoretical justification for the method of ERM. This does not exclude that one might (also) wish for different kinds of justification, in any specific problem or in general. In any case, my aim here is to flag the above qualifications, to make clear that accepting the justification for ERM, and indeed the justification for simplicity to follow, presupposes accepting those qualifications. One can reject the argument below simply by rejecting the presuppositions of the statistical learning theory framework. For instance, the reasons Kelly (2008, 2011) offers for rejecting arguments for Occam’s razor from statistical learning theory are essentially a rejection of the predictive framework. [^28] Kelly (2008, p. 329) writes that Occam’s razor “should help one to select the true theory from among the alternatives,” whereas arguments based on risk minimization do not concern “theoretical truth” but “passive prediction” (ibid., 335). “But beliefs are for guiding action and actions can alter the world so that the sampling distribution we drew our conclusions from is altered as well” (ibid.); moreover, “it is clear that the over-fitting story depends, essentially, upon noise in the data […] One would prefer that the connection between simplicity and theoretical truth not depend essentially upon randomness” (ibid.). These points are all well-taken (in particular the problem of distribution-shift has recently received more attention, Wiles et al., 2022), but are all already concerns about the scope of statistical learning theory itself. But, I will argue, conditional on those qualifications—including the presuppositions of the framework—we can formulate a justificatory argument.
@@
There is, however, a final aspect to this justification that requires more discussion. Namely, it can be seen to cover only one side of an inevitable trade-off.
@@ 4.1.2. The bias-complexity trade-off
The opposing pull of these two error terms is also referred to as the bias-complexity trade-off. A lower complexity—lower capacity—class excludes more possibilities, and as such embodies, in machine learning terminology, a stronger inductive bias. [^29] This is a bit more general than the bias-variance trade-off (see Hastie et al., 2009, sects. 2.6, 2.9). Complexity and (inductive) bias are here in the first instance used as informal terms (even if complexity can be made precise as VC dimension, and some authors refer to the approximation error itself as the bias), while the bias and variance in the latter are well-defined statistical terms in regression with mean squared error. The fundamental theorem yields a guarantee about finding the best in a given class, but this is inevitably a class with some inductive bias. The resulting justification must therefore be relative to this class or inductive bias. [^30] Of course, the hypothesis class is not the only—or even the most important (von Luxburg and Schölkopf, 2011, p. 684)—way in which assumptions or biases enter: as discussed in section 4.1 above, important modeling assumptions must already be made in the formalization of the learning problem (including choice of feature space and loss function). But discussions of inductive bias (in particular around the no-free-lunch theorems introduced shortly) usually assume that these elements are already in place, and center on the further inductive assumptions required.
@@
In the terminology of Sterkenburg and Grünwald (2021), this is a model-relative justification. It is indeed a justification for a learning method, ERM, that is explicitly model-dependent. The method of ERM is a general procedure—a “generic learning rule” (Shalev-Shwartz and Ben-David, 2014, p. 68)—yet one that must, by definition, on each application be supplied with further assumptions. We can view ERM as instantiating a two-place function, that apart from a data sample, also takes a particular hypothesis class. On each specific application, it must be supplied with a hypothesis class or inductive model that constitutes further (context-dependent) assumptions, the inductive bias. Correspondingly, the learnability guarantee for ERM is model-relative, because the notion of learnability is relative to a given inductive model or VC hypothesis class.
@@
The analysis of Sterkenburg and Grünwald aims to explain how general learning-theoretic guarantees for generic algorithms are consistent with the skeptical import of the so-called no-free-lunch theorems of supervised learning (going back to Wolpert, 1992, 1996; Schaffer, 1994). Modern versions of these results (Shalev-Shwartz and Ben-David, 2014, pp. 36ff; Sterkenburg and Grünwald, 2021, pp. 9990f) say that there can exist no universal learning algorithm: every particular algorithm is inadequate in some possible learning situations, situations where another algorithm is adequate. And since there can be no a priori justification for privileging particular learning situations, so the further interpretation goes, there can be no theoretical justification for any particular algorithm.
@@
However, rather than the generic yet model-dependent ERM algorithm, the no-free-lunch statement applies to any particular instantiation of ERM with an inductive model $\mathcal{H}$, any particular one-place “data-only” function ERMH. The no-free-lunch result of Shalev-Shwartz and Ben-David (2014, thrm. 5.1) essentially states that for any specific inductive model $\mathcal{H}$, the data-only algorithm ERMH is inadequate (i.e., with high probability suffers high absolute error) for some situations (i.e., for some true distributions; informally, those that do not match $\mathcal{H}$’s inductive bias) where another data-only algorithm, like ${}_{\mathcal{H^{\prime}}}$ for another inductive model $\mathcal{H}^{\prime}$ (that does match the situation), is adequate. [^31] Another way of casting this result is that the class of all classifiers is not learnable: since, for any possible distribution, this class has minimum approximation error, its learnability (guarantee of low estimation error) would guarantee low absolute error for any possible distribution. In fact, the proof of Shalev-Shwartz and Ben-David (2014, thrm. 5.1) already shows the failure of learnability of classes with infinite VC dimension, and the no-free-lunch theorem is in their presentation part of the proof of the fundamental theorem (ibid., pp. 45ff).
@@
In other words, while we can have a model-relative justification for model-dependent algorithms (of the rough form, “for any instantiated inductive model of the right form, works well relative to the model”), there is no absolute justification (“works well whenever”) for any inductive model. The impossibility of such absolute justification is also an important part of the argument of Domingos (1998, 1999) against a possible justification of Occam’s razor.
@@ 4.1.4. The failure of an absolute justification for simplicity
Domingos takes issue with what he calls the “second razor”: that “given two [hypotheses] with the same [empirical] error, the simpler one should be preferred because it is likely to have lower generalization error” (1999, p. 410). (To be distinguished from the “first razor,” that “the simpler [hypothesis] should be preferred because simplicity is desirable in itself,” ibid.) [^32] We are here, of course, likewise concerned with an epistemic justification for Occam’s razor, not with claims that simplicity is better because it points to hypotheses or theories that are easier to work with or more aesthetically pleasing (cf. Sober, 2015, pp. 58f). He writes that theoretical “zero-sum arguments”—no-free-lunch theorems—“imply that the second razor cannot be true” (1999, p. 413). Namely, “they imply that, for every domain where a simpler [hypothesis] is more accurate than a more complex one, there exists a domain where the reverse is true, and thus no argument which is preferable in general can be made” (ibid.).
This is true, and an expression of the impossibility of absolute justification, applied to the choice of a simple single hypothesis. [^33] This point has also been brought out experimentally (e.g., Schaffer, 1993; Webb, 1996). However, I already observed that the relevant notion of simplicity attaches to hypothesis classes, not single hypotheses. Domingos’s critique of the theoretical “PAC-learning argument” for the second razor, which is also a no-free-lunch observation, is more relevant:
@@
4.2. Epistemology and machine learning theory
4.2.1. “Model-relative” epistemology